Trains A & B are traveling in the same direction on parallel tracks. Tran A is traveling 80 miles per hour and Train B is traveling 90 miles per hour. Train A passes a station at 12:00am. If Train B passes the same station 12:40am, at what time will Train B catch up to Train A? When will Train B catch up with Train A? What is the time and if am or pm? Please help. Thank you.
Answer by mananth(16946)

Train A speed = 80 mph
Train B 90 mph
Train A is ahead of train B at the station by 40 minutes.( 2/3 hours)
By the time train B reaches the station train A has traveled 80 * 2/3 miles
=160/3 miles to a point say P.
Let the distance from point P to the point where train B overtakes train A be x
.
Train B travels 160/3 + x miles in the same time the train A travels x miles
time taken by train A = x/80 hours
time taken by train B = x+160/3 / 90
.
x/80 = 3x+160 / 3 divided by 90
x/80 = 3x+160 /3 * 1/90
x/80= 3x+ 160 / 270
270x = 80(3x+160)
270x=240x +12800
30x= 12800
x=426.66 miles
time taken by train A t travel x miles = 426.66/80
=5.33 hours.
This is the same time taken by train B to cross train A